Prove that every field is artinian.Deduce that Q is artinian.


1
Expert's answer
2021-06-02T16:49:26-0400

Let us prove that every field is Artinian. Recall that an Artinian ring is a ring that satisfies the descending chain condition on ideals, i.e. for any decreasing sequence of right ideals "I_1\supset I_2\supset\ldots" there exists a natural number m such that "I_m=I_{m+1}=\ldots" .


Taking into account that each field "F" contains exactly two ideals: "F" and "O=\{0\}", we conclude that for each descending chain of ideals there are two possibilites:

1) Each ideal in a chain is equal to "F".

2) There exists an ideal "I_m" that is equal to "O". Since we consider only descending chains, "I_k=O" for every integer "k>m".


We conclude that in each case a field satisfies the descending chain condition on ideals, and hence each field is Artinian.


Since "(\mathbb Q,+,\cdot)" is a field, it is Artinian.